Question Help
In a club with

7 male and

11 female​ members, a

3​-member committee will be randomly chosen. Find the probability that the committee contains at least

2 women.
1
Expert's answer
2019-04-04T10:00:35-0400

We have a committee that will have 3 people and at least 2 of them will be women. There are 7 men and 11 women that can be on the committee. How many ways can we select 3 people?


"C(18, 3)=\begin{pmatrix}
 18 \\
 3
\end{pmatrix}={18! \over 3!(18-3)!}={18(17)(16) \over 1(2)(3)}=816"

There are "\begin{pmatrix}
 7 \\
 3
\end{pmatrix}" ways to choose three men out of seven and "\begin{pmatrix}
 11 \\
 0
\end{pmatrix}" ways to choose zero women out of eleven.


"\begin{pmatrix}
 7 \\
 3
\end{pmatrix}\begin{pmatrix}
 11 \\
 0
\end{pmatrix}={7! \over 3!(7-3)!}*{11! \over 0!(11-0)!}=35"

There are"\begin{pmatrix}
 7 \\
 2
\end{pmatrix}" ways to choose two men out of seven and"\begin{pmatrix}
 11 \\
 1
\end{pmatrix}" ways to choose one woman out of eleven.


"\begin{pmatrix}
 7 \\
 2
\end{pmatrix}\begin{pmatrix}
 11 \\
 1
\end{pmatrix}={7! \over 2!(7-2)!}*{11! \over 1!(11-1)!}=231"

Find the probability that the committee contains at least 2 women.


"P(W \geq 2)=1-{35+231\over816}={275 \over 408}\approx0.674"